OpenStudy (anonymous):

Write the equation of the line that passes through (1, 5) and (−2, 14) in slope-intercept form.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

A. y = 3x + 2 B. y = 3x + 8 C. y = −3x − 2 D. y = −3x + 8

jimthompson5910 (jim_thompson5910):

what is the change in y when we go from y = 5 to y = 14 ?

OpenStudy (anonymous):

I have no idea.

jimthompson5910 (jim_thompson5910):

subtract the values to find out

OpenStudy (anonymous):

9?

jimthompson5910 (jim_thompson5910):

yes, now what is the change in x when we go from x = 1 to x = -2 ?

OpenStudy (anonymous):

-3?

jimthompson5910 (jim_thompson5910):

yes -2-1 = -3

jimthompson5910 (jim_thompson5910):

change in y = 9 change in x = -3 divide the two values: 9 divided by -3 = -3 so the slope here is -3

OpenStudy (anonymous):

So, it's either C or D?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

the slope is m = -3 we replace m with -3 in y = mx+b so y = mx+b turns into y = -3x+b

jimthompson5910 (jim_thompson5910):

now pick one of the two points, say (1,5) the first coordinate is x, which means x = 1 the second coordinate is y, so y = 5 you plug x = 1 and y = 5 into y = -3x+b to get y = -3x+b 5 = -3*1+b now isolate b

OpenStudy (anonymous):

So, you just have, 5=-3*1?

jimthompson5910 (jim_thompson5910):

5 = -3*1+b turns into 5 = -3+b which is the same as 5 = b-3

jimthompson5910 (jim_thompson5910):

you have to solve for b now

OpenStudy (anonymous):

How do you do that?

jimthompson5910 (jim_thompson5910):

how would you undo that -3 ?

OpenStudy (anonymous):

Dividing?

jimthompson5910 (jim_thompson5910):

addition is the opposite of subtraction

OpenStudy (anonymous):

So we add?

OpenStudy (anonymous):

So, would it be, D. y = −3x + 8?

jimthompson5910 (jim_thompson5910):

yes add 3 to both sides and you end up with b = 8 so the final answer is y = -3x+8

OpenStudy (anonymous):

Thank you!!

jimthompson5910 (jim_thompson5910):

no problem